# A remark on $\Lambda^{2}$-enlargeable manifolds

### v1

@@ Abstract.
In this note, we consider the case that the condition “constant near infinity” in the definition of $\Lambda^{2}$-enlargeable manifold replaced by the condition “locally constant near infinity” and prove that $\Lambda^{2}$-enlargeable manifold in the current sense also can not carry a complete Riemannian metric of positive scalar curvature. As a consequence, we give another proof of Wang-Zhang’s theorem on the generalized Geroch conjecture for complete spin manifolds.
@@ 1. Introduction
$\Lambda^{2}$-enlargeable manifold was introduced by Gromov-Lawson ([2]). A famous theorem of Gromov-Lawson ([2]) states that $\Lambda^{2}$-enlargeable manifold can not carry a complete Riemannian metric of positive scalar curvature.
Let $W$ be a closed $\Lambda^{2}$-enlargeable manifold and $M$ be noncompact connected spin manifold without boundary with ${\rm dim}M={\rm dim}W$, Wang and Zhang ([8]) proved that the connected sum M#WM\#W can not carry a complete Riemannian metric of positive scalar curvature using the result in [9]. If $M$ is a closed spin manifold, then M#WM\#W is a closed $\Lambda^{2}$-enlargeable manifold ([2]). So it is natural to ask that whether M#WM\#W is a $\Lambda^{2}$-enlargeable manifold for the case that $M$ is a noncompact spin manifold. For noncompact $M$, from the constructions in [7, 8] (which goes back to [2]), one finds that the condition that the maps from the covering manifolds to the standard sphere are constant near infinity does not satisfy for M#WM\#W. In fact the maps are locally constant near infinity. In this note, we consider the $\Lambda^{2}$-enlargeable manifold in this case and prove that the $\Lambda^{2}$-enlargeable manifold in the current sense also can not carry a complete Riemannian metric of positive scalar curvature.
@@
In the following definition of $\Lambda^{2}$-enlargeable Riemannian metric, we replace “constant near infinity” in [2, Definition 7.1] by “locally constant near infinity”. In [5, Definition 1.10], Shi also considered this case.
@@
A Riemannian metric on a connected manifold $M$ is called $\Lambda^{2}$-enlargeable if given any $\epsilon>0$, there exists a covering manifold $M_{\epsilon}\to M$ such that $M_{\epsilon}$ is spin and a smooth map $f_{\epsilon}:M_{\epsilon}\to S^{{\rm dim}M}(1)$ which is $(\epsilon,\Lambda^{2})$-contracting with respect to the lifted metric, locally constant near infinity and of non-zero degree. A connected (not necessarily compact) manifold is said to be $\Lambda^{2}$-enlargeable if any Riemannian metric (not necessarily complete) on $M$ is $\Lambda^{2}$-enlargeable.
@@
A proof of Theorem 1.3 will be given in Section 2.
Lemma 1.4.
Proof.
Let $g^{T(M\#W)}$ be any Riemannian metric on T(M#W)T(M\#W). We fix a point $p\in W$. For any $r\geq 0$, let $B^{W}_{p}(r)=\{y\in W:d(p,y)\leq r\}$. Let $b_{0}>0$ be a fixed sufficiently small number. Then the connected sum M#WM\#W can be constructed so that the hypersurface $\partial B^{W}_{p}(b_{0})$, which is the boundary of $B^{W}_{p}(b_{0})$, cuts M#WM\#W into two parts: the part $W\setminus B^{W}_{p}(b_{0})$ and the rest part coming from $M$ (by attaching the boundary of a ball in $M$ to $\partial B^{W}_{p}(b_{0})$).
From the metric $g^{T(M\#W)}$, by restriction we get a metric $g^{T(W\setminus B^{W}_{p}(b_{0}))}$ on $T(W\setminus B^{W}_{p}(b_{0}))$, and we extend it to a smooth metric $g^{TW}$ on $TW$. Then $g^{TW}$ is a $\Lambda^{2}$-enlargeable metric. For any $\epsilon>0$, let $\pi:\widehat{W}_{\epsilon}\to W$ be a covering manifold verifying [2, Definition 7.1], carrying lifted geometric data from that of $W$. Especially there exists a smooth map $f:\widehat{W}_{\epsilon}\to S^{{\rm dim}M}(1)$ and a compact subset $K_{\epsilon}\subset\widehat{W}_{\epsilon}$ such that $f$ is constant on $\widehat{W}_{\epsilon}\setminus K_{\epsilon}$. As in [7, 8], the connected sum M#WM\#W lifts naturally to $\widehat{W}_{\epsilon}$ and we denote the resulting manifold by $\widehat{M}\#\widehat{W}_{\epsilon}$. We lift the metric $g^{T(M\#W)}$ to $\widehat{M}\#\widehat{W}_{\epsilon}$. We extend the map $f$ to $\widehat{M}\#\widehat{W}_{\epsilon}$ by seeting that $f|_{\widehat{M}\#(\widehat{W}_{\epsilon}\setminus K_{\epsilon})}$ is constant. Then by the construction in [8], we have a map $\widehat{f}:\widehat{M}\#\widehat{W}_{\epsilon}\to S^{{\rm dim}M}(1)$ which is $(c\epsilon,\Lambda^{2})$-contracting with respect to the lifted metric for some constant $c>0$, locally constant near infinity and of non-zero degree. Then the metric $g^{T(M\#W)}$ is a $\Lambda^{2}$-enlargeable metric in the sense of Definition 1.2 and by definition M#WM\#W is a $\Lambda^{2}$-enlargeable manifold in the sense of Definition 1.2.
By Theorem 1.3 and Lemma 1.4, we get [8, Theorem 1.1], which states that M#WM\#W can not carry a complete Riemannian metric of positive scalar curvature.
2. A proof of Theorem 1.3
@@
Following the proof of [2, Theorem 6.12], we consider another metric on $TM$ defined by $k^{TM}g^{TM}$. By definition, for the metric $k^{TM}g^{TM}$ and any $\epsilon>0$, there exists a covering
@@
such that $M_{\epsilon}$ is spin and a smooth map
@@
For $M_{\epsilon}$, $g^{TM_{\epsilon}}$ and $f_{\epsilon}:M_{\epsilon}\to S^{n}(1)$, we do the same constructions as in [4, Section 3.1] and we use the same notations except that we add $\epsilon$ to denote the manifold $M_{\epsilon}$. Then we have a compact manifold $M_{H_{\epsilon,3m}}$ with boundary $H_{\epsilon,3m}$ and a closed manifold $\widehat{M}_{H_{\epsilon,3m}}$ with the metric $g_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}}$. Let $\nabla_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}}$ be the associated Levi-Civita connection. The map $f_{\epsilon}|_{M_{H_{\epsilon,3m}}}$ also extends to a map $f_{\epsilon,l}:\widehat{M}_{H_{\epsilon,3m}}\to S^{n}(1)$.
We first assume that $n$ is even.
@@
Let $S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})=S_{\beta,+}(T\widehat{M}_{H_{\epsilon,3m}})\oplus S_{\beta,-}(T\widehat{M}_{H_{\epsilon,3m}})$ be the $\mathbb{Z}_{2}$-graded Hermitian vector bundle of spinors associated to $(T\widehat{M}_{H_{\epsilon,3m}},g_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}})$, carrying the canonical induced Hermitian connection $\nabla^{S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})}=\nabla^{S_{\beta,+}(T\widehat{M}_{H_{\epsilon,3m}})}\oplus\nabla^{S_{\beta,-}(T\widehat{M}_{H_{\epsilon,3m}})}$.
Let $S(TS^{n}(1))=S_{+}(TS^{n}(1))\oplus S_{-}(TS^{n}(1))$ be the spinor bundle of $S^{n}(1)$. Let $V:S(TS^{n}(1))\to S(TS^{n}(1))$ be the map defined in [6, Section 2.5]. Then there exists $\delta>0$ such that
@@
Let $D^{E_{3m}}_{\beta}$ acting on $S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})\widehat{\otimes}E_{3m}$ be the twisted Dirac operators defined by
@@
(2.9) $\left|\left({1\over 2}\sum_{i,j=1}^{n}R^{E_{3m,r}}(h_{i},h_{j})c_{\beta}(h_{i})c_{\beta}(h_{j})s,s\right)(x)\right|\\ =\left|\left({1\over 2}\sum_{i,j}^{n}f^{*}_{\epsilon,l,r}(R^{S}(f_{\epsilon,l,r,*}(h_{i}\wedge h_{j})))c_{\beta}(h_{i})c_{\beta}(h_{j})s,s\right)(x)\right|\leq{{\epsilon k^{TM_{\epsilon}}}\over{2\beta^{2}}}n(n-1)C_{1}|s|^{2}(x),$
@@
Then using (2.1), (2.6), (2.9) and (2.11)-(2.14), proceeding as in the proof of [6, Lemma 2.1(i)] and [6, Section 2.5], one finds that there exist $c_{0}>0,\varepsilon>0,m>0$ such that when $\beta>0$ is small enough,
@@
On the other hand, by the Atiyah-Singer index theorem [1] (cf. [3, Proposition III. 13.8]), as in [6, (2.44)], we have
(2.16) ${\rm ind}\left(\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)_{+}\right)=2(-1)^{n\over 2}{\rm deg}(f_{\epsilon})\neq 0.$
@@
If $n$ is odd, we consider the composition $f_{\epsilon,l,r}$ of the maps

### v2

@@ Abstract.
In this note, we consider the case where the condition “constant near infinity” in the definition of $\Lambda^{2}$-enlargeable manifolds is replaced by the condition “locally constant near infinity” and prove that a $\Lambda^{2}$-enlargeable manifold in this modified sense still cannot carry a complete Riemannian metric of positive scalar curvature. As a consequence, we give another proof of Wang-Zhang’s theorem on the generalized Geroch conjecture for complete spin manifolds.
@@ 1. Introduction
The $\Lambda^{2}$-enlargeable manifold was introduced by Gromov-Lawson ([4]). A famous theorem of Gromov-Lawson ([4]) states that a $\Lambda^{2}$-enlargeable manifold cannot carry a complete Riemannian metric of positive scalar curvature.
Let $W$ be a closed $\Lambda^{2}$-enlargeable manifold and $M$ be a noncompact connected spin manifold without boundary with ${\rm dim}M={\rm dim}W$. Wang and Zhang ([10]) proved that the connected sum M#WM\#W cannot carry a complete Riemannian metric of positive scalar curvature using the result in [11]. If $M$ is a closed spin manifold, then M#WM\#W is a closed $\Lambda^{2}$-enlargeable manifold ([4]). So it is natural to ask whether M#WM\#W is a $\Lambda^{2}$-enlargeable manifold for the case that $M$ is a noncompact spin manifold. For a noncompact $M$, from the constructions in [9, 10] (which goes back to [4]), one finds that the condition that the maps from the covering manifolds to the standard sphere are constant near infinity is not satisfied for M#WM\#W. In fact, the maps are only locally constant near infinity. In this note, we consider the $\Lambda^{2}$-enlargeable manifold in this case and prove that the $\Lambda^{2}$-enlargeable manifold in the current sense also cannot carry a complete Riemannian metric of positive scalar curvature.
@@
In the following definition of $\Lambda^{2}$-enlargeable Riemannian metrics, we replace “constant near infinity” in [4, Definition 7.1] by “locally constant near infinity”. In [7, Definition 1.10], Shi also considered this condition, but from a different motivation, also, there is no overlap between Shi’s main results and the results proved in the present paper.
@@
Let $M$ be a connected manifold without boundary. A Riemannian metric on $M$ is called $\Lambda^{2}$-enlargeable if given any $\epsilon>0$, there exist a covering manifold $M_{\epsilon}\to M$ such that $M_{\epsilon}$ is spin and a smooth map $f_{\epsilon}:M_{\epsilon}\to S^{{\rm dim}M}(1)$ which is $(\epsilon,\Lambda^{2})$-contracting with respect to the lifted metric, locally constant near infinity and of non-zero degree. A connected (not necessarily compact) manifold is said to be $\Lambda^{2}$-enlargeable if all the Riemannian metrics (not necessarily complete) on $M$ are $\Lambda^{2}$-enlargeable.
Proposition 1.3.
From Proposition 1.3 one finds that there are manifolds which are $\Lambda^{2}$-enlargeable in the sense of Definition 1.2 but not $\Lambda^{2}$-enlargeable in the sense of [4, Definition 7.1]. For example, since the torus $T^{n}$ is a closed $\Lambda^{2}$-enlargeable manifold, then $M\#T^{n}$ is a $\Lambda^{2}$-enlargeable manifold in the current sense for a noncompact spin manifold $M$.
The main theorem of this paper can be stated as follows, which extends [4, Theorem 6.12] to the current case.
@@
By Proposition 1.3 and Theorem 1.4, we give another proof of [10, Theorem 1.1], which states that M#WM\#W cannot carry a complete Riemannian metric of positive scalar curvature.
The rest of this paper is organized as follows. In Section 2, we will give a proof of Proposition 1.3. In Section 3, we will give a proof of Theorem 1.4.
2. Proof of Proposition 1.3
Let $g^{TW}_{0}$ be a fixed Riemannian metric on $TW$. We fix a point $p\in W$. For any $r\geq 0$, let $B^{W}_{p}(r)=\{y\in W:d(p,y)\leq r\}$. As in [9, 10], let $b_{0}$ be a fixed sufficiently small number. Then the connected sum M#WM\#W can be constructed so that the hypersurface $\partial B^{W}_{p}(b_{0})$, which is the boundary of $B^{W}_{p}(b_{0})$, cuts M#WM\#W into two parts: one part $W\setminus B^{W}_{p}(b_{0})$ and the other part coming from $M$ (by attaching the boundary of a ball in $M$ to $\partial B^{W}_{p}(b_{0})$). Let $g_{0}^{T(W\setminus B^{W}_{p}(b_{0}))}$ be the restricted metric on $T(W\setminus B^{W}_{p}(b_{0}))$ from $g^{TW}_{0}$.
Let $g^{T(M\#W)}$ be an arbitrary Riemannian metric on T(M#W)T(M\#W). From the metric $g^{T(M\#W)}$, by restriction we get a metric $g^{T(W\setminus B^{W}_{p}(b_{0}))}$ on $T(W\setminus B^{W}_{p}(b_{0}))$. For any $\epsilon>0$, let $\pi:\widehat{W}_{\epsilon}\to W$ be a covering manifold satisfying [4, Definition 7.1], carrying the lifted geometric data of $(W,g^{TW}_{0})$. Especially there exist a smooth map $f:\widehat{W}_{\epsilon}\to S^{{\rm dim}M}(1)$ and a compact subset $K_{\epsilon}\subset\widehat{W}_{\epsilon}$ such that $f$ is constant on $\widehat{W}_{\epsilon}\setminus K_{\epsilon}$. As in [9, 10], the connected sum M#WM\#W is lifted naturally to $\widehat{W}_{\epsilon}$ and we denote the resulting manifold by $\widehat{M}\#\widehat{W}_{\epsilon}$. We lift the metric $g^{T(M\#W)}$ to $\widehat{M}\#\widehat{W}_{\epsilon}$ and extend the map $f$ to $\widehat{M}\#\widehat{W}_{\epsilon}$ by setting that $f|_{\widehat{M}\#(\widehat{W}_{\epsilon}\setminus K_{\epsilon})}$ is constant. Since $\overline{W\setminus B^{W}_{p}(b_{0})}$ is compact, the metrics $g^{T(W\setminus B^{W}_{p}(b_{0}))}_{0}$ and $g^{T(W\setminus B^{W}_{p}(b_{0}))}$ are equivalent. Then by the construction in [10], we have a map $\widehat{f}:\widehat{M}\#\widehat{W}_{\epsilon}\to S^{{\rm dim}M}(1)$ which is $(c\epsilon,\Lambda^{2})$-contracting with respect to the lifted metric of $g^{T(M\#W)}$ for some constant $c>0$, locally constant near infinity and of non-zero degree. Then the metric $g^{T(M\#W)}$ is a $\Lambda^{2}$-enlargeable metric in the sense of Definition 1.2 and by definition M#WM\#W is a $\Lambda^{2}$-enlargeable manifold in the sense of Definition 1.2.
3. Proof of Theorem 1.4
In this section we give a proof of Theorem 1.4 using the methods in [8] and [6].
@@
Following the proof of [4, Theorem 6.12], we consider another metric on $TM$ defined by $k^{TM}g^{TM}$. By definition, for the metric $k^{TM}g^{TM}$ and any $\epsilon>0$, there exists a covering
@@
such that $M_{\epsilon}$ is spin and there exists a smooth map
@@
3.1. Construct a closed manifold
In this subsection, we briefly recall from [8] and [6] the construction of a closed manifold from $(M_{\epsilon},g^{TM_{\epsilon}})$.
Following [4, Theorem 1.17], we choose a fixed point $x_{0}\in M_{\epsilon}$ and let $d:M_{\epsilon}\to\mathbb{R}^{+}$ be a regularization of the distance function ${\rm dist}(x,x_{0})$ such that
(3.1) $|\nabla d|(x)\leq\frac{3}{2},$
for any $x\in M_{\epsilon}$. Set
(3.2) $B_{\epsilon,m}=\{x\in M_{\epsilon}:d(x)\leq m\},\ m\in\mathbb{N}.$
Since $K_{\epsilon}$ is compact, we can choose a sufficiently large $m$ such that $K_{\epsilon}\subseteq B_{\epsilon,m}$. This implies
(3.3) ${\rm Supp}({\rm d}f_{\epsilon})\subseteq K_{\epsilon}\subseteq B_{\epsilon,m}.$
Following [4], we take a compact hypersurface $H_{\epsilon,3m}\subseteq M_{\epsilon}\setminus K_{\epsilon}$, cutting $M_{\epsilon}$ into two parts such that the compact part, denoted by $M_{H_{\epsilon,3m}}$, contains $B_{\epsilon,3m}$. Then $M_{H_{\epsilon,3m}}$ is a compact smooth manifold with boundary $H_{\epsilon,3m}$. Note that the number of connected components of $M_{\epsilon}\setminus M_{H_{\epsilon,3m}}$ is finite. Let $\{Y_{j}\}_{j=1}^{l}$ be the connected components of $M_{\epsilon}\setminus M_{H_{\epsilon,3m}}$.
Let $H_{\epsilon,3m}\times[-1,2]$ be the product manifold and we construct a metric $H_{\epsilon,3m}\times[-1,2]$ as in [8].
Assume $f(Y_{j})=p_{j}\in S^{n}(1)$, $j=1,\dots,l$. We choose a point $p_{0}\in S^{n}(1)$ and for $j=1,\dots,l$, pick a curve $\xi_{j}(\tau),0\leq\tau\leq 1$, connecting $p_{j}$ and $p_{0}$ such that $\xi_{i}(\tau)\cap\xi_{j}(\tau^{\prime})=\emptyset,0<\tau,\tau^{\prime}<1,i\neq j$. Following [8], for $(y,\tau)\in H_{\epsilon,3m}\times[-1,2]$ and $j=1,\dots,l$, we define [^1] A similar trick also appears in [3].
(3.4) $f_{\epsilon}(y,\tau)=\begin{cases}p_{j},\ (y,\tau)\in(Y_{j}\cap H_{\epsilon,3m})\times[-1,0],\\ \xi_{j}(\tau),\ (y,\tau)\in(Y_{j}\cap H_{\epsilon,3m})\times[0,1],\\ p_{0},\ (y,\tau)\in(Y_{j}\cap H_{\epsilon,3m})\times[1,2].\end{cases}$
Note that some points of $\{p_{j}\}_{j=1}^{l}$ may coincide. Without loss of generality, we assume that $(0,\cdots,0,\pm 1)\notin\xi_{j}(\tau),1\leq j\leq l$. The map $f_{\epsilon}$ can be extended to a map on $M_{H_{\epsilon,3m}}\cup(H_{\epsilon,3m}\times[-1,2])$ via $f_{\epsilon}(y,\tau)$. Denote such a map on $M_{H_{\epsilon,3m}}\cup(H_{\epsilon,3m}\times[-1,2])$ by $f_{\epsilon,l}$.
Let $M^{\prime}_{H_{\epsilon,3m}}$ be another copy of $M_{H_{\epsilon,3m}}$ with the same metric and the opposite orientation. As in [8], we can glue $M_{H_{\epsilon,3m}}$, $H_{\epsilon,3m}\times[-1,2]$ and $M^{\prime}_{H_{\epsilon,3m}}$ together to get a closed manifold $\widehat{M}_{H_{\epsilon,3m}}$. We view $M_{H_{\epsilon,3m}}$, $M^{\prime}_{H_{\epsilon,3m}}$ and $H_{\epsilon,3m}\times[-1,2]$ as submanifolds of $\widehat{M}_{H_{\epsilon,3m}}$. The map $f_{\epsilon,l}$ can be extended to $\widehat{M}_{H_{\epsilon,3m}}$ by setting $f_{\epsilon,l}(M^{\prime}_{H_{\epsilon,3m}})={p_{0}}$. We still denote the map on $\widehat{M}_{H_{\epsilon,3m}}$ by $f_{\epsilon,l}$. The map $f_{\epsilon,l}$ has the following properties:
(3.5) ${\rm Supp}({\rm d}f_{\epsilon,l})\subseteq{\rm Supp}({\rm d}f_{\epsilon})\cup(H_{\epsilon,3m}\times[0,1]),\ {\rm deg}(f_{\epsilon,l})={\rm deg}(f_{\epsilon})\neq 0.$
For any $\beta>0$, let $g^{TM_{\epsilon}}_{\beta}$ be the Riemannian metric on $M_{\epsilon}$ defined by
(3.6) $g^{TM_{\epsilon}}_{\beta}=\beta^{2}g^{TM_{\epsilon}}.$
Let $g^{TH_{\epsilon,3m}}$ be the induced metric on $H_{\epsilon,3m}$ by (3.6) with $\beta=1$ and ${\rm d}t^{2}$ be the standard metric on $[0,1]$. By the construction of $\widehat{M}_{H_{\epsilon,3m}}$, we can define a smooth metric $g_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}}$ on $\widehat{M}_{H_{\epsilon,3m}}$ in the following way:
(3.7) $\left.g_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}}\right|_{M_{H_{\epsilon,3m}}}=g^{TM_{\epsilon}}_{\beta},\ \left.g_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}}\right|_{M^{\prime}_{H_{\epsilon,3m}}}=g^{TM_{H_{\epsilon,3m}}^{\prime}},\ \left.g_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}}\right|_{H_{\epsilon,3m}\times[0,1]}=g^{TH_{\epsilon,3m}}\oplus{{\rm d}t^{2}},$
and then paste these metrics together. Let $\nabla_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}}$ be the Levi-Civita connection on $T\widehat{M}_{H_{\epsilon,3m}}$ associated with the metric $g_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}}$.
3.2. The even-dimensional case
@@
In this subsection, we assume $n$ is even. Let $S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})=S_{\beta,+}(T\widehat{M}_{H_{\epsilon,3m}})\oplus S_{\beta,-}(T\widehat{M}_{H_{\epsilon,3m}})$ be the $\mathbb{Z}_{2}$-graded Hermitian vector bundle of spinors associated with $(T\widehat{M}_{H_{\epsilon,3m}},g_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}})$, carrying the canonical induced Hermitian connection $\nabla^{S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})}=\nabla^{S_{\beta,+}(T\widehat{M}_{H_{\epsilon,3m}})}\oplus\nabla^{S_{\beta,-}(T\widehat{M}_{H_{\epsilon,3m}})}$.
Let $S(TS^{n}(1))=S_{+}(TS^{n}(1))\oplus S_{-}(TS^{n}(1))$ be the spinor bundle of $S^{n}(1)$. Following [11, (2.6)], we construct a suitable bundle endomorphism $V$ of $S(TS^{n}(1))$. More precisely, by taking any regular value $\mathfrak{q}\in S^{n}(1)\setminus f_{\epsilon}(M_{\epsilon}\setminus K_{\epsilon})$ of $f_{\epsilon}$, we choose $X$ to be a smooth vector field on $S^{n}(1)$ such that $|X|>0$ on $S^{n}(1)\setminus\{\mathfrak{q}\}$. Let
$v=c(X):S_{+}(TS^{n}(1))\to S_{-}(TS^{n}(1))$
be the Clifford action of $X$ and
$v^{*}:S_{-}(TS^{n}(1))\to S_{+}(TS^{n}(1))$
be the adjoint of $v$ with respect to the Hermitian metric on $S_{\pm}(TS^{n}(1))$. We define $V$ to be the self-adjoint odd endomorphism
$V=v+v^{*}:S(TS^{n}(1))\to S(TS^{n}(1)).$
Then there exists $\delta>0$ such that
@@
Let $D^{E_{3m}}_{\beta}$ acting on $S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})\widehat{\otimes}E_{3m}$ be the twisted Dirac operator defined by
@@
(3.16) $\left|\left({1\over 2}\sum_{i,j=1}^{n}R^{E_{3m}}(h_{i},h_{j})c_{\beta}(h_{i})c_{\beta}(h_{j})s,s\right)(x)\right|\\ =\left|\left({1\over 2}\sum_{i,j}^{n}f^{*}_{\epsilon,l}(R^{S}(f_{\epsilon,l,*}(h_{i}\wedge h_{j})))c_{\beta}(h_{i})c_{\beta}(h_{j})s,s\right)(x)\right|\leq{{\epsilon k^{TM_{\epsilon}}}\over{2\beta^{2}}}n(n-1)C_{1}|s|^{2}(x),$
@@
Following [4, Theorem 1.17], let $\phi:[0,\infty)\rightarrow[0,1]$ be a smooth function such that $\phi\equiv 1$ on $[0,1]$, $\phi\equiv 0$ on $2,\infty)$ and $\phi^{\prime}\approx-1$ on $[1,2]$. We define a smooth function ${\psi_{m}}:M_{H_{\epsilon,3m}}\to[0,1]$ by
(3.22) $\psi_{m}(x)=\phi\left(\frac{d(x)}{m}\right),$
where $m\in\mathbb{N}$. We extend $\psi_{m}$ to $(H_{\epsilon,3m}\times[-1,2])\cup M^{\prime}_{H_{\epsilon,3m}}$ by setting
$\psi_{m}\big((H_{\epsilon,3m}\times[-1,2])\cup M^{\prime}_{H_{\epsilon,3m}}\big)=0.$
Following [2, p. 115], let $\psi_{m,1},\,\psi_{m,2}:\widehat{M}_{H_{\epsilon,3m}}\rightarrow[0,1]$ be defined by
(3.23) $\psi_{m,1}=\frac{\psi_{m}}{\big(\psi_{m}^{2}+(1-\psi_{m})^{2}\big)^{\frac{1}{2}}},\;\psi_{m,2}=\frac{1-\psi_{m}}{\big(\psi_{m}^{2}+(1-\psi_{m})^{2}\big)^{\frac{1}{2}}}.$
Using the above definition and (3.1), for $i=1,2$, we have
(3.24) $|\nabla\psi_{m,i}|(x)\leq{C/m}\text{ for any }x\in\widehat{M}_{H_{\epsilon,3m}},$
where $C$ is a constant independent of $g^{TM_{\epsilon}}$.
For any $s\in\Gamma(\widehat{M}_{H_{\epsilon,3m}},S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})\widehat{\otimes}E_{3m})$, by (3.23), one has
(3.25) $\left\|\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)s\right\|^{2}_{\beta}\\ =\left\|\psi_{m,1}\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)s\right\|^{2}_{\beta}+\left\|\psi_{m,2}\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)s\right\|^{2}_{\beta},$
from which one gets
(3.26) $\sqrt{2}\left\|\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)s\right\|_{\beta}\\ \geq\left\|\psi_{m,1}\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)s\right\|_{\beta}+\left\|\psi_{m,2}\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)s\right\|_{\beta}\\ \geq\left\|\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)(\psi_{m,1}s)\right\|_{\beta}+\left\|\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)(\psi_{m,2}s)\right\|_{\beta}\\ -\|c_{\beta}({\rm d}\psi_{m,1}s)\|_{\beta}-\|c_{\beta}({\rm d}\psi_{m,2}s)\|_{\beta},$
where for each $i\in\{1,2\}$, we identify ${\rm d}\psi_{m,i}$ with the gradient of $\psi_{m,i}$.
For any $x\in M_{H_{\epsilon,3m}}$ and $i\in\{1,2\}$, as in [8, (2.19)], we have
(3.27) $|c_{\beta}({\rm d}\psi_{m,i})s|_{\beta}(x)=O\left({1\over{\beta m}}\right)|s|_{\beta}(x).$
Now we estimate the first term on the right-hand side of (3.26). Using (3.8), (3.13), (3.16)-(3.18) and (3.21), we have
(3.28) $\left\|\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)(\psi_{m,1}s)\right\|^{2}_{\beta}\\ \geq\left({{k^{TM_{\epsilon}}}\over{4\beta^{2}}}\psi_{m,1}s,\psi_{m,1}s\right)_{\beta}+\left({1\over 2}\sum_{i,j=1}^{n}R^{E_{3m}}(h_{i},h_{j})c_{\beta}(h_{i})c_{\beta}(h_{j})\psi_{m,1}s,\psi_{m,1}s\right)_{\beta}\\ +\left(\left[D^{E_{3m}}_{\beta},{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right]\psi_{m,1}s,\psi_{m,1}s\right)_{\beta}+\left(\left({{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)^{2}\psi_{m,1}s,\psi_{m,1}s\right)_{\beta}\\ \geq{\kappa\over{8\beta^{2}}}\|s\|^{2}_{\beta,{\rm Supp}({\rm d}f_{\epsilon})}+O_{\epsilon}\left({\varepsilon\over{\beta^{2}}}\right)\|s\|^{2}_{\beta,{\rm Supp}({\rm d}f_{\epsilon})}+{{\delta\varepsilon^{2}}\over{\beta^{2}}}\|\psi_{m,1}s\|^{2}_{\beta,B_{2m}\setminus{\rm Supp}({\rm d}f_{\epsilon})}.$
For the second term on the right-hand side of (3.26), using (3.8) and (3.20), we have
(3.29) $\left\|\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)(\psi_{m,2}s)\right\|^{2}_{\beta}\geq{{\delta\varepsilon^{2}}\over{\beta^{2}}}\|\psi_{m,2}s\|^{2}_{\beta}+O_{m,\epsilon}\left({\varepsilon\over\beta}\right)\|\psi_{m,2}s\|^{2}_{\beta,H_{\epsilon,3m}\times[0,1]}.$
Then using (3.26)-(3.29), one finds that there exist $c_{0}>0,\varepsilon>0,m>0$ such that when $\beta>0$ is small enough,
@@
On the other hand, by the Atiyah-Singer index theorem [1] (cf. [5, Proposition III. 13.8]), as in [8, (2.44)], we have
(3.31) ${\rm ind}\left(\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)_{+}\right)\\ =\left\langle\widehat{{A}}(TM_{\epsilon})f_{\epsilon,l}^{*}\left({\rm ch}(S_{+}(TS^{n}(1)))-{\rm ch}(S_{-}(TS^{n}(1)))\right),[M_{\epsilon}]\right\rangle\\ ={\rm deg}(f_{\epsilon,l})\Big\langle{\rm ch}(S_{+}(TS^{n}(1)))-{\rm ch}(S_{-}(TS^{n}(1))),[S^{n}(1)]\Big\rangle\\ =(-1)^{n\over 2}{\rm deg}(f_{\epsilon})\chi(S^{n}(1))=2(-1)^{n\over 2}{\rm deg}(f_{\epsilon})\neq 0.\\$
@@
3.3. The odd-dimensional case
In this subsection, we assume $n$ is odd. We consider the composition $f_{\epsilon,l,r}$ of the maps